OpenStudy (anonymous):

Add and write in simplest form: 4y/y+2 + 8/y+2

OpenStudy (anonymous):

4(y+2)/(y+2)=4

OpenStudy (anonymous):

4y+8=4(y+2) and as both have the same denominator the common factor is y+2 so answer is 4